Parth (parthkohli):

How to add fractions like these? \[\LARGE {2 \over x + 1} + {3 \over x + 2}\]

OpenStudy (anonymous):

you can cross multiply and get 3(x+1)=2(x+2)

OpenStudy (anonymous):

oh like these okay so We have x+1 • 3 = 3x+3 x+2 • 2= 2x + 2 then just add that *like a boss*

jhonyy9 (jhonyy9):

not is right sure

Parth (parthkohli):

Actually, 2(x + 2) = 2x + 4. reb

OpenStudy (anonymous):

hahaha omg my bad...ooppsie

Parth (parthkohli):

Can you explain why?

OpenStudy (anonymous):

why what?

Parth (parthkohli):

Why they are cross multiplied?

jhonyy9 (jhonyy9):

2/(x+1) + 3/(x+2) =(2(x+2)+3(x+1))/(x+1)(x+2) = =(2x+4+3x+3)/(x+1)(x+2) =(5x+7)/(x+1)(x+2)

jhonyy9 (jhonyy9):

so because (x+1)(x+2) is the denominator common

OpenStudy (anonymous):

@jhonyy9 is absolutely correct

OpenStudy (anonymous):

you can cross multiply to find x 3(x+1) 3x+3=3x/3=1 2(x+2) 2x+4=2x/2=2

Parth (parthkohli):

Oh! I get it....

jhonyy9 (jhonyy9):

hope that this was understandably sure

Parth (parthkohli):

You made me clear! Thanks!!!

jhonyy9 (jhonyy9):

ok yw good luck bye